# softMathematical software

### v1

@@ 2.2.1. Discrete probabilistic CI models
2.2.2. Semi-graphoids
@@
The concept of a semi-graphoid was introduced by Pearl [26] with the aim to formalize fundamental properties of probabilistic CI pinpointed earlier by Dawid [9]. Here we give an equivalent definition from [20, § 5] adapted to our situation.
A semi-graphoid over a variable set $N$ is a subset ${\mathcal{M}}$ of ${\mathcal{K}}(N)$ satisfying
@@
The frame $\mathfrak{sg}$ is closed under copying, marginalization, intersection, and duality.
Proof 2.7.
This basically follows from the above definition of a semi-graphoid.
Well-known facts are that $\mathfrak{pr}\subseteq\mathfrak{sg}$ and $\mathfrak{sg}(N)=\mathfrak{pr}(N)$ if $|N|\leq 3$. The simplest example that $\mathfrak{sg}(N)\neq\mathfrak{pr}(N)$ in case $|N|\geq 4$ is as follows.
@@
Given $N:=\{a,b,c,d\}$ and :=${\mathcal{M}}:=\{\,ab|c,\,ac|d,\,ad|b\,\}\in\mathfrak{sg}(N)$, one has ${\mathcal{M}}\not\in\mathfrak{pr}(N)$ by property (E:1) in [31, § V.A] saying that the equivalence
@@
Definition 2.9 (global CI statement).
We interpret this as a global CI statement encoded within the model ${\mathcal{M}}$. In our examples and catalogues we also use a shorthand $[I,J|K]$ to denote the list of (elementary) CI statements described in (1).
Despite the global CI statement is formally defined relative to any CI model ${\mathcal{M}}$, the concept itself makes sense only when ${\mathcal{M}}$ is a semi-graphoid. This relates to two alternative modes for describing semi-graphoids: in terms of either global or local (= elementary) CI statements; see [20, § 5] or [29, Lemma  2.2].
Remark. This is to elucidate possible alternative ways to formalize CI models. An equivalent way to the one from Definition 2.1 is to replace ${\mathcal{K}}(N)$ by the collection ${\mathcal{T}}(N)$ of all triplets $\langle I,J|K\rangle$ of pairwise disjoint subsets $I,J,K$ of $N$ and define (global) CI models as subsets of ${\mathcal{T}}(N)$. These two equivalent approaches are then related through (the concept introduced in) Definition 2.9.
A non-equivalent approach is to replace ${\mathcal{K}}(N)$ by the collection ${\mathcal{K}}^{\bullet}(N)$ of ordered pairs $\,L\,|\,K\,$, where $L,K\subseteq N$, $L\cap K=\emptyset$, and $1\leq|L|\leq 2$. Any pair $i\,|\,K$ then corresponds to a functional dependence statement claiming that $i\in N\setminus K$ is “determined” by $K\subseteq N$, interpreted as a specific CI statement $i\perp\!\!\!\perp i\,|\,K$ (= $i$ is conditionally independent of itself given $K$). The subsets of ${\mathcal{K}}^{\bullet}(N)$ are then augmented CI models over $N$ [31, § II.B] involving functional dependencies, represented in the local/elementary mode. This wider approach can equivalently be described by replacing ${\mathcal{K}}^{\bullet}(N)$ with the collection ${\mathcal{T}}^{\bullet}(N)$ of all triplets $\langle I,J|K\rangle$ of subsets $I,J,K$ of $N$. The subsets of ${\mathcal{T}}^{\bullet}(N)$ can be viewed as augmented CI models, represented in the global mode. However, we do not consider augmented CI models with functional dependence statements in this paper.
2.2.3. Structural semi-graphoids
The concept of a structural semi-graphoid was introduced in [28] as a semi-graphoid induced in a particular way by certain subsets of ${\mathbb{Z}}^{{\mathcal{P}}(N)}$. The original complicated definition was shown in [29, Corollary 5.3] to be equivalent to a much simpler condition that underlies our definition below.

### v3

@@ 2.1. Abstract families of CI models and the probabilistic CI frame
Remark. This is to elucidate possible alternative ways to formalize CI models. The classic way of understanding probabilistic CI structures from [32], followed in [37], was to define an abstract CI model over $N$ as the collection of global CI statements over $N$ valid with respect to a random vector $\xi$ over $N$. This approach leads to an alternative global mode, where ${\mathcal{K}}(N)$ is replaced by the collection ${\mathcal{T}}(N)$ of all triplets $\langle I,J|K\rangle$ of pairwise disjoint subsets $I,J,K$ of $N$ and (global) CI models are subsets of ${\mathcal{T}}(N)$. Such an alternative approach is equivalent to the one discussed in our paper through the relation from Definition 2.3.
[…]
2.2. Tutorial on Galois connections and formal concept analysis
This part contains an overview of particular results from lattice theory and formal concept analysis we later apply in our context of abstract CI frames (see Section 2.3) to establish frame-based implications of CI models. We have included such a summary in our paper because a common reader need not be familiar with these finer results, perhaps well-known in lattice-theoretical community, but not in general. The reason is that the consistency of our algorithms (Appendix A) follows just from these facts.
The main message of this tutorial to be passed towards the reader is that any finite lattice, possibly with a huge number of elements, can (efficiently) be characterized in two methodologically different ways:
  • in terms of irreducible elements (= an inner description in the sense that the characterization only depends on the lattice itself),
Below we describe both approaches and recall relevant concepts from lattice theory [4] for this purpose.
Throughout this tutorial, $X$ will be a general set. Facts from lattice theory hold for any such a set $X$. Finer concepts, which concern the implications between subsets of $X$, and results from formal concept analysis implicitly assume that $X$ is finite. We are going to apply those observations in Section 2.3 in the case $X:={\mathcal{K}}(N)$ for some finite variable set $N$.
An elementary observation is that any Moore family $({\mathcal{F}},\subseteq)$, equipped with the set inclusion relation, is a complete lattice [4, Theorem 2 in § V.1] in which the meet operation is intersection. [^2] This is a universal construction of a complete lattice in the sense that any such a lattice is isomorphic to a Moore family and any finite lattice is isomorphic to a finite Moore family lattice. These conclusions can be derived from observations in [4, § V.9]. Hence, the following simple facts can be derived.
Observation 2.7.
If a CI frame $\mathfrak{f}$ is closed under intersection then, for each finite $N$, the family $(\mathfrak{f}(N),\subseteq)$ is a finite lattice in which the meet operation is intersection. If $\mathfrak{f}$ is also closed under copying and $|N|=|M|$ then $(\mathfrak{f}(N),\subseteq)$ and $(\mathfrak{f}(M),\subseteq)$ are order-isomorphic and permutations of $N$ yield automorphisms of $(\mathfrak{f}(N),\subseteq)$.
Proof.
By Definition 2.4, as $\mathfrak{f}(N)$ is a Moore family, this is evident from what was said before. ∎
Let $({\mathcal{L}},\preceq)$ be a (finite) lattice with the greatest element denoted by 1 and the meet/infimum operator denoted by $\inf$. Let $\prec$  denote the strict order relation. Given $\lambda,\tau\in{\mathcal{L}}$, one says that $\tau$ covers $\lambda$ and writes $\lambda\prec\!\!\!\cdot\,\,\,\tau$ if $\lambda\prec\tau$ and there is no $\kappa\in{\mathcal{L}}$ with $\lambda\prec\kappa\prec\tau\/$. A coatom of ${\mathcal{L}}$ is a sub-maximal element in ${\mathcal{L}}$, that is, $\tau\in{\mathcal{L}}$ satisfying $\tau\prec\!\!\!\cdot\,\,\,\mbox{\bf 1}$. An element $\lambda\in{\mathcal{L}}$ is called (meet) irreducible if $\lambda\neq\inf\,\{\,\tau\in{\mathcal{L}}\,:\ \lambda\prec\tau\,\}$. [^3] Note that 1 is not irreducible by this definition: it is the infimum of the empty set. A well-known fact is that the set ${\mathcal{I}}$ of (meet) irreducible elements in a finite lattice $({\mathcal{L}},\preceq)$ is the least set determining all elements of ${\mathcal{L}}$ by means of the infimum operation in the following sense:
The coatoms are standard examples of meet irreducible elements in ${\mathcal{L}}$. In general, a lattice can have other irreducible elements. A lattice $({\mathcal{L}},\preceq)$ is called coatomistic if every element of ${\mathcal{L}}$ is the infimum of a set of coatoms, which is another way of saying that there is no irreducible element in ${\mathcal{L}}$ except the coatoms.
Example 2.8.
Put $X:=\{x,y,z,w\}$ and consider a Moore family
:=${\mathcal{F}}:=\{\,\,\{x,y,z,w\},\{x,y,z\},\{x,y,w\},\{x,y\},\{x\}\,\,\}\,,$
Figure 1. A Moore family $\{\,\{x,y,z,w\},\{x,y,z\},\{x,y,w\},\{x,y\},\{x\}\,\}$.
The first way to characterize a finite lattice $({\mathcal{F}},\subseteq)$, where ${\mathcal{F}}$ is a Moore family of subsets of $X$, is to give the complete list of irreducible elements in ${\mathcal{F}}$. In our considered case they fall into equivalence classes:
Observation 2.9.
If a CI frame $\mathfrak{f}$ is closed under copying and intersection then, for each finite variable set $N$, the meet irreducible sets in $(\mathfrak{f}(N),\subseteq)$ (= irreducible models) fall into permutational equivalence classes (= relative to permutations of $N$).
Proof.
By Observation 2.7, any permutation $\pi:N\to N$ defines an order automorphism of $(\mathfrak{f}(N),\subseteq)$, the natural extension of $\pi$ to ${\mathcal{K}}(N)$. Such a transformation preserves the irreducibility of an element of $\mathfrak{f}(N)$. ∎
Example 2.10.
(lattice of probabilistic models over 3 variables)
Let us take $N:=\{a,b,c\}$, $X:={\mathcal{K}}(N)$, and put ${\mathcal{F}}:=\mathfrak{pr}\,(N)$. Either a human-made analysis or simple computations allows one to observe that the lattice $(\mathfrak{pr}\,(N),\subseteq)$ has 22 elements in this case and they break into 10 permutational equivalence classes; for the result see [37, Figure 5.6]. There are 5 (meet) irreducible elements in $\mathfrak{pr}\,(\{a,b,c\})$ which fall down into 3 permutational types:
  • $\{\,ab|\emptyset,\,ac|\emptyset,\,bc|\emptyset\,\}$,
  • $\{\,ab|c,\,ac|b,\,bc|a\,\}$,
A kind of dual description of a complete lattice is based on the following concept. A closure operation on subsets of $X$ is a mapping $\mbox{\rm cl}:{\mathcal{P}}(X)\to{\mathcal{P}}(X)$ which is
  • extensive: $Y\subseteq\mbox{\rm cl}\,(Y)$ for $Y\subseteq X$,
A set $Y\subseteq X$ is then called closed with respect to cl if $Y=\mbox{\rm cl}\,(Y)$. A basic facts are that the collection of subsets of $X$ closed with respect to cl is a Moore family and any Moore family can be defined in this way [4, Theorem 1 in § V.1]. Specifically, given a Moore family ${\mathcal{F}}$ of subsets of $X$, the formula
defines a closure operation on subsets of $X$ having ${\mathcal{F}}$ as the collection of closed sets with respect to $\mbox{\rm cl}_{{\mathcal{F}}}$; see [18, Theorem 1 in § 0.3] for the proof of this simple observation. Thus, for any $Y\subseteq X$, the closure $\mbox{\rm cl}_{{\mathcal{F}}}(Y)$ of $Y$ relative to ${\mathcal{F}}$ is the smallest element of ${\mathcal{F}}$ containing $Y$.
Example 2.11.
The closure operation can also be interpreted as an abstract functional dependence relation on subsets of $X$, as explained in [24]. This is a binary relation $\to$ on ${\mathcal{P}}(X)$ satisfying three axioms (for $Y,Z,W\subseteq X$):
  • reflexivity: $Z\subseteq Y~{}\Rightarrow~{}Y\to Z$,
  • transitivity: $[\,Y\to Z~{}\&~{}Z\to W\,]~{}\Rightarrow~{}Y\to W$,
  • union: $[\,Y\to Z~{}\&~{}Y\to W\,]~{}\Rightarrow~{}Y\to Z\cup W$.
In case of finite $X$, these axioms are together equivalent to the well-known Armstrong axioms for functional dependence relations in database theory. Given a closure operation cl on subsets of a finite set $X$, define $Y\to Z$ as $Z\subseteq\mbox{\rm cl}\,(Y)$ and observe it is an abstract functional dependence relation on subsets of $X$. Given a functional dependence relation $\to$, put $\mbox{\rm cl}\,(Y):=\bigcup\,\{\,Z\subseteq X\,:\ Y\to Z\,\}$ for $Y\subseteq X$ and observe this defines a closure operation on subsets of $X$ inducing the original functional dependence relation $\to$.
Example 2.12.
[…]
Remark. To avoid possible terminological misunderstanding on the side of the reader note that functional dependencies within a random vector $\xi$ mentioned in the remark concluding Section 2.1 can be viewed as a part of an abstract probabilistic functional dependence relation. Specifically, given a random vector $\xi$ over a finite variable set $N$, the collection of pairs $K\to I$, where $K,I\subseteq N$, $K\cap I=\emptyset$, and $I\perp\!\!\!\perp I\,|\,K\,\,[\mbox{\boldmath$\xi$}]$ (augmented global CI statements) is an abstract functional dependence relation on subsets of $X:=N$. One of the results in [24, Remark  4] is that every abstract functional dependence relation on subsets of a finite $N$ is probabilistic in the above sense. So, these notions are indeed related.
Nonetheless, in our paper we ignore probabilistic functional dependencies within $N$. Instead, we deal with functional dependence on a meta-level of CI statements/models over $N$: we consider abstract functional dependence relations on subsets of $X:={\mathcal{K}}(N)$, where $N$ is a finite variable set. To emphasize the different context, we also talk about “implications" between subsets of $X$, which are CI models over $N$.
The second way to characterize a finite lattice $({\mathcal{F}},\subseteq)$, where ${\mathcal{F}}$ is a Moore family of subsets of $X$, is to give a suitable generator of the respective functional dependence relation $\to$. Before explaining the details we wish the reader to realize/observe that there is a straightforward relationship (= correspondence) between the functional dependence relations $\to$ and Moore families ${\mathcal{F}}$ [^4] A one-to-one assignment between Moore families and functional dependence relations has already been established through their shared closure operations. We now re-interpret this as a kind of duality relationship and avoid the closure operations.
Specifically, any element $(Y,Z)$ of ${\mathcal{P}}(X)\times{\mathcal{P}}(X)$, where $X$ is a finite set, can be interpreted as a formal implication $Y\to Z$, meaning that the set $Y$ of antecedents implies (all elements of) the set $Z$ of consequents. One can introduce a particular binary relation $\odot$ between such pairs $(Y,Z)$ and subsets $F$ of $X$: we say that $Y\to Z$ is valid for $F\subseteq X$ if $Y\subseteq F$ implies $Z\subseteq F$. More technically:
$(Y\to Z)\odot F\quad:=\quad[\,\,Y\setminus F\neq\emptyset~{}~{}\mbox{or}~{}~{}Z\subseteq F\,\,].$
Next we introduce the so-called Galois connections between subsets of ${\mathcal{P}}(X)\times{\mathcal{P}}(X)$ and subsets of ${\mathcal{P}}(X)$:
Well-known classic results in lattice theory on Galois connections from [4, § V.7], completed by particular observations from [24, (unnumbered) Theorem], say the following in case of a finite set $X$:
  • the mapping ${\mathbb{G}}\mapsto{\mathbb{G}}^{\triangleright\triangleleft}$ is a closure operation on subsets of the set ${\mathcal{P}}(X)\times{\mathcal{P}}(X)$ and ${\mathbb{G}}^{\triangleright\triangleleft}$ is the smallest functional dependence relation (on subsets of $X$) containing ${\mathbb{G}}$,
  • the mapping ${\mathcal{F}}\mapsto{{\mathcal{F}}}^{\triangleleft\triangleright}$ is a closure operation on subsets of ${\mathcal{P}}(X)$ and ${{\mathcal{F}}}^{\triangleleft\triangleright}$ is the smallest Moore family (of subsets of $X$) containing ${\mathcal{F}}$,
Example 2.13.
Consider $X:=\{x,y,z,w\}$ and put
${\mathbb{G}}:=\{~{}\emptyset\to\{x\},~{}\{z\}\to\{y\},~{}\{x,w\}\to\{x,y,w\}~{}\}\,.$
Then ${\mathbb{G}}^{\triangleright}$ equals to the family ${\mathcal{F}}$ from Example 2.8. Hence, ${\mathbb{G}}$ is one of several possible implicational generators of ${\mathcal{F}}$, a minimal one in sense of inclusion (= a non-redundant one).
To introduce a prominent implicational generator one needs specific concepts from the theory of formal concept analysis; see [16, Chapter  3] or [40, § 1.3.2]. [^5] In general, there is no unique smallest implicational generator is sense of inclusion, that is, a non-redundant generator.
Definition 2.14 (pseudo-closed sets, canonical basis).
Given a finite set $X$, a subset $Q\subseteq X$ is pseudo-closed relative to a Moore family ${\mathcal{F}}\subseteq{\mathcal{P}}(X)$ if
  • $Q\not\in{\mathcal{F}}$ (= $Q$ is not closed),
The recursive definition of pseudo-closedness of a set $Q$ is well-formed as it depends only on the previously defined pseudo-closedness of strict subsets of $Q$. The canonical implicational basis, alternatively named Guigues–Duquenne basis, was introduced already in the 1980s [15]. We have also implemented an algorithm to compute the canonical basis ${\mathbb{G}}_{\mathcal{F}}$ for a given Moore family ${\mathcal{F}}$; its idea is described in Appendix A.
Example 2.15.
$\{~{}\emptyset\to\{x\},\,\{z\}\to\{y\},\,\{w\}\to\{y\}~{}\}\,.$
The widely-known result in formal concept analysis says that $({\mathbb{G}}_{{\mathcal{F}}})^{\triangleright}={\mathcal{F}}$ for any Moore family ${\mathcal{F}}\subseteq{\mathcal{P}}(X)$; see [16, Theorem 7 in § 3.2] for a relatively short proof. In fact, ${\mathbb{G}}_{{\mathcal{F}}}$ is a non-redundant implicational generator of ${\mathcal{F}}$ in sense that the removal of any of its elements leads to a non-generator.
Let us introduce further specific terminology concerning formal implications. Given a Moore family ${\mathcal{F}}\subseteq{\mathcal{P}}(X)$, a valid implication for ${\mathcal{F}}$ is $(Y\to Z)\in{\mathcal{F}}^{\triangleleft}$. An implication $Y\to Z$ is perfect (for ${\mathcal{F}}$) if $Y\not\in{\mathcal{F}}$ but every strict subset of $Y$ is closed, that is, $Y\in({\mathcal{P}}(X)\setminus{\mathcal{F}})_{\min}$, and $Z=\mbox{\rm cl}_{{\mathcal{F}}}(Y)\setminus Y$. The implication is canonical if it belongs to the canonical implicational basis ${\mathbb{G}}_{{\mathcal{F}}}$.
Definition 2.16 (implicatively perfect Moore family).
It makes no problem to deduce from non-redundancy of the canonical basis that an equivalent condition is that every pseudo-closed set is an inclusion-minimal non-closed set, that is, an element of set collection $({\mathcal{P}}(X)\setminus{\mathcal{F}})_{\min}$. Let us illustrate the above defined notions by an example.
Example 2.17.
(probabilistic models over 3 variables revisited)
Consider again the situation from Example 2.10: $N:=\{a,b,c\}$, $X:={\mathcal{K}}(N)$, and ${\mathcal{F}}:=\mathfrak{pr}\,(N)$. There are six pseudo-closed sets relative to ${\mathcal{F}}$ belonging to one permutational type: the set $\{\,ab|\emptyset,\,ac|b\,\}$ and five of its permutational copies. Each of these sets is minimal in ${\mathcal{P}}(X)\setminus{\mathcal{F}}$; thus, ${\mathcal{F}}$ is implicatively perfect. The canonical basis ${\mathbb{G}}_{{\mathcal{F}}}$ consists of permutational copies of the implication
$\{\,ab|\emptyset,\,ac|b\,\}\to\{\,ac|\emptyset,\,ab|c\,\}\,.$
The characterization of this particular Moore family ${\mathcal{F}}$ in terms of its canonical implicational basis is equivalent to a later definition of a semi-graphoid over $\{a,b,c\}$ in Section 3.1.
Remark. The concept from Definition 2.16 can be viewed as a counter-part of the one of a coatomistic lattice in the context of implicational description of the lattice. Realize that the Galois connection ${\mathcal{F}}\mapsto{\mathcal{F}}^{\triangleleft}$, for Moore families ${\mathcal{F}}$ of subsets of a finite set $X$, can be viewed as an anti-isomorphism between such Moore families and abstract functional dependence relations ${\mathbb{G}}$, whose inverse is the Galois connection ${\mathbb{G}}\mapsto{\mathbb{G}}^{\triangleright}$. We can interpret this map as a kind of self-inverse duality transformation between these mathematical structures, which both can be viewed as alternative (ways of) descriptions of finite lattices. If one prefers the perspective of a Moore family then the simplest lattice is the coatomistic one but if one prefers the perspective of implicational description then the simplest lattice is the one described by perfect implications.
2.3. Frame-based closure of a CI model and implications between CI models
Let us come back to the situation in Section 2.1 and consider an abstract CI frame $\mathfrak{f}$ closed under intersection. In this case, for every finite variable set $N$, the family of CI models $(\mathfrak{f}(N),\subseteq)$ is a Moore family of subsets of $X:={\mathcal{K}}(N)$. Therefore, the closure operation $\mbox{\rm cl}_{\mathfrak{f}(N)}$ on subsets of ${\mathcal{K}}(N)$, respectively the collection of valid implications for $\mathfrak{f}(N)$, characterizes this family. Observe that, for ${\mathcal{M}},{\mathcal{R}}\subseteq{\mathcal{K}}(N)$, the validity of the implication ${\mathcal{M}}\to{\mathcal{R}}$ for $\mathfrak{f}(N)$ reduces to the condition
$\forall\,{\mathcal{E}}\in\mathfrak{f}(N)\quad{\mathcal{M}}\subseteq{\mathcal{E}}~{}\mbox{implies}~{}{\mathcal{R}}\subseteq{\mathcal{E}}\,.$
In the sequel, we intend to switch freely between the description of a CI frame $\mathfrak{f}$ in terms of the families $\mathfrak{f}(N)$ and in terms of the corresponding closure operations.
The fact that the elements of the canonical basis in Example 2.17 fall down into permutational types is not a surprise because we have the following.
Observation 2.18.
If a CI frame $\mathfrak{f}$ is closed under copying and intersection then, for each set $N$, the valid implications ${\mathcal{M}}\to{\mathcal{R}}$ for the Moore family ${\mathcal{F}}:=\mathfrak{f}(N)$ and the elements of the canonical implicational basis for $\mathfrak{f}(N)$ fall into permutational equivalence classes.
Proof.
The arguments are analogous to those in Observation 2.9. Any permutation $\pi:N\to N$ defines a simultaneous order automorphism of the posets $(\mathfrak{f}(N),\subseteq)$ and $({\mathcal{P}}({\mathcal{K}}(N)),\subseteq)$. Because such an order automorphism of $({\mathcal{F}},\subseteq)$ commutes with the closure operation $\mbox{\rm cl}_{{\mathcal{F}}}$, it preserves the validity of implications. ∎
We typically wish to have the frame-based closure stable relative to an enlargement of the basic set.
Observation 2.19.
Let $\mathfrak{f}$ be a CI frame closed under marginalization, intersection, and ascetic extension. Given ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ and $N\subseteq M$ one has $\mbox{\rm cl}_{\mathfrak{f}(N)}({\mathcal{M}})=\mbox{\rm cl}_{\mathfrak{f}(M)}({\mathcal{M}})$.
Proof.
The fact that $\mathfrak{f}$ is closed under marginalization implies $\mbox{\rm cl}_{\mathfrak{f}(N)}({\mathcal{M}})\subseteq\mbox{\rm cl}_{\mathfrak{f}(M)}({\mathcal{M}})$. Since $\mathfrak{f}(N)\subseteq\mathfrak{f}(M)$ (ascetic extension), one has $\mbox{\rm cl}_{\mathfrak{f}(N)}({\mathcal{M}})\in\mathfrak{f}(M)$, which yields the equality. ∎
Therefore, the assumption on a CI frame $\mathfrak{f}$ being closed under ascetic extension makes the following definition of the (frame-based) $\mathfrak{f}$-closure of a CI model ${\mathcal{M}}$ consistent, because the closure does not depend on the “context", by which we mean a variable set $M$ such that ${\mathcal{M}}\subseteq{\mathcal{K}}(M)$.
Definition 2.20 (frame-based closure of a CI model).
Consider a CI frame $\mathfrak{f}$ closed under marginalization, intersection, and ascetic extension. Given a CI model ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ over a variable set $N$, its closure of relative to $\mathfrak{f}(N)$ will be denoted by $\mathfrak{f}({\mathcal{M}}):=\mbox{\rm cl}_{\mathfrak{f}(N)}({\mathcal{M}})$. We will call it the $\mathfrak{f}$-closure of the model ${\mathcal{M}}$.
Indeed, by Observation 2.19, if ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$ then both ${\mathcal{M}}$ and $\mathfrak{f}({\mathcal{M}})$ are models over $N^{{\mathcal{M}}}:=\bigcup_{ij|K\in{\mathcal{M}}}ijK$. Also, if the assumptions of Definition 2.20 are fulfilled and ${\mathcal{M}}\subseteq{\mathcal{K}}(N)$, ${\mathcal{R}}\subseteq{\mathcal{K}}(M)$ are given then the condition ${\mathcal{M}}\to{\mathcal{R}}$ for $\mathfrak{f}(NM)$ enforces $N^{{\mathcal{R}}}\subseteq N^{{\mathcal{M}}}$ and its verification reduces to showing that ${\mathcal{R}}\subseteq\mathfrak{f}({\mathcal{M}})$.
3. Examples of abstract CI frames
This is an overview of CI frames; some of these were topics of our computational experiments.
3.1. Semi-graphoids
@@
The concept of a semi-graphoid was introduced by Pearl [32] with the aim to formalize fundamental properties of probabilistic CI pinpointed earlier by Dawid [10]. Here we give an equivalent definition from [25, § 5] adapted to our situation. A semi-graphoid over a variable set $N$ is a subset ${\mathcal{M}}$ of ${\mathcal{K}}(N)$ satisfying
@@
$\mathfrak{sg}$ is closed under copying, marginalization, intersection, duality, and ascetic extension.
Proof.
This basically follows from the above definition of a semi-graphoid. ∎
Well-known facts are that $\mathfrak{pr}\subseteq\mathfrak{sg}$ and $\mathfrak{sg}(N)=\mathfrak{pr}(N)$ if $|N|\leq 3$. The simplest example showing that $\mathfrak{sg}(N)\neq\mathfrak{pr}(N)$ in case $|N|\geq 4$ is as follows.
@@
Given $N:=\{a,b,c,d\}$ and :=${\mathcal{M}}:=\{\,ab|c,\,ac|d,\,ad|b\,\}\in\mathfrak{sg}(N)$, one has ${\mathcal{M}}\not\in\mathfrak{pr}(N)$ by property (E:1) in [39, § V.A] saying that the equivalence
@@
3.2. Structural semi-graphoids
The concept of a structural semi-graphoid was introduced in [36] as a semi-graphoid induced in a particular way by certain subsets of ${\mathbb{Z}}^{{\mathcal{P}}(N)}$. The original complicated definition was shown in [37, Corollary 5.3] to be equivalent to a much simpler condition that underlies our definition below.